# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ IV-C DoF of the Cascade Link
$C=\sum\limits_{k=1}^{K}\log_{2}\left(1+\left(\tilde{\sigma}_{k}^{(\mathrm{T})}\right)^{2}\left(\tilde{\sigma}_{k}^{(\mathrm{R})}\right)^{2}\frac{p_{k}}{\sigma_{\mathrm{noise}}^{2}}\right)$ (35)
@@
where $p_{k}=\max\left(\mu-\frac{\sigma_{\mathrm{noise}}^{2}}{\left(\tilde{\sigma}_{k}^{(\mathrm{T})}\right)^{2}\left(\tilde{\sigma}_{k}^{(\mathrm{R})}\right)^{2}},0\right)$, with $\mu$ obtained by fulfilling the identity $\sum\nolimits_{k=1}^{K}{p_{k}}=P_{\text{T}}$, and being $\tilde{\sigma}_{k}^{(\mathrm{T})}$ and $\tilde{\sigma}_{k}^{(\mathrm{R})}$ the singular values of ${\bf{H}}_{\mathrm{T}}$ and ${\bf{H}}_{\mathrm{R}}$, respectively. [^6] We have $\tilde{\sigma}_{k}^{(\mathrm{T})}=\sigma_{k}^{(\mathrm{T})}$, since ${\bf{V}}$ is unitary. In contrast, $\tilde{\sigma}_{k}^{(\mathrm{R})}$ may differ from $\sigma_{k}^{(\mathrm{R})}$ due to the presence of the matrix ${\bf{G}}^{-1}$ in ${\bf{H}}_{\mathrm{R}}$, which accounts for the coupling between the current density and the scattered electric field at the reconfigurable scattering device. It is important to note that the last $K-R$ power coefficients are zero due to the maximum rank of $\mathbf{H}(\bm{\theta})$ being $R$. Specifically, because of the diagonal structure of $\mathbf{G}_{\text{T}}$ and $\mathbf{G}_{\text{R}}$, where only the first $N$ and $M$ diagonal entries, respectively, correspond to significant singular values, only the first $R$ elements of the symbol vector ${\bf{x}}$ (out of $K$) carry “usable” information data. In addition, (35) is obtained by setting $\mathbf{Q}={\rm diag}\left(p_{1},p_{2},\ldots p_{K}\right)$ and the coupling matrix ${\bf{C}}={\bf{I}}_{K}$. Under the Born approximation, from (19), the optimal device matrix should satisfy the condition
${\bf{R}}(\bm{\theta})={\bf{V}}_{\text{R}}{\bf{U}}^{{\mathrm{H}}}_{\text{T}}$ (36)
@@
where ${\bf{V}}_{\text{R}}$ and ${\bf{U}}_{\text{T}}$ represent the left and right eigenvector matrices of the SVD of ${\bf{H}}_{\mathrm{T}}$ and ${\bf{H}}_{\mathrm{R}}$, respectively. From (36), it can be inferred that the end-to-end capacity is maximized by designing ${\bf{R}}(\bm{\theta})$ (and consequently ${\bf{D}}(\bm{\theta})$) in a way that aligns with the left eigenvectors of the first link and the right eigenvectors of the second link. This ensures that the end-to-end transmitter-device-receiver channel is diagonalized. Furthermore, the water-filling power allocation is applied to the ordered product of the singular values of the individual transmitter-device and device-receiver channels.
@@
The physical realizability of the optimal device matrix $\mathbf{R}(\bm{\theta})$ depends on the specific technology being used. As mentioned in Sec. III, when the boundary conditions imposed by the scattering device are local, such as in conventional RISs, the matrix $\mathbf{R}(\bm{\theta})$ is constrained to be diagonal, meaning that (36) can only be approximated [44]. Recently, new RIS structures have been introduced, which enable reconfigurable coupling between the elements, leading to a non-diagonal matrix (non-diagonal RISs) [45].
V Enabling EM Technologies and SP Challenges
@@
In this section, we provide a comprehensive overview of key EM enabling technologies that can be utilized within the ESP framework by presenting specific examples of their application in practical processing tasks.
@@
V-A Port-controlled Reconfigurable Scattering Devices
@@
In most practical implementations, the reconfigurable scattering device is composed of a discrete set of $K$ antenna elements located at positions ${\bf{r}}_{k}$ whose scattering characteristics can be adjusted by changing the load at their port. With a limited loss of generality, for a given $k$-th antenna element of the reconfigurable scattering device, define ${\bf{F}}_{k}({\bf{r}})$ $\,[1/m^{2}]$ the corresponding current density normalized to 111\,1[A] current at its input port. Here we are implicitly assuming that the antenna element can be characterized by only one dominant mode which is a good approximation, for instance, for Hertizian dipoles and resonant structures such as the half-wave dipole. The function ${\bf{F}}_{k}({\bf{r}})$ depends on the particular structure of the element as well as its position ${\bf{r}}_{k}$ and orientation. Such a distribution can be obtained in general through EM-based simulations, but closed-form expressions are available for simple structures such as Hertzian or half-wave dipoles [35]. The entire scattering device can be considered as a linear $K$-port network where we define the complex voltage and current envelopes (in the following denoted simply as voltages and currents) at the ports, respectively, ${\bf{v}}$ (open circuit voltage) and ${\bf{i}}_{\mathrm{s}}$. The actual current density of the $k$-th element is therefore ${{\mathsf{J}}_{\mathrm{s}}}_{k}({\bf{r}})=[{{\bf{i}}_{\mathrm{s}}}]_{k}\cdot{\bf{F}}_{k}({\bf{r}})$. The corresponding unitary-energy basis function is ${\mathbf{\Phi}}_{k}({\bf{r}})={\bf{F}}_{k}({\bf{r}})/\|{\bf{F}}_{k}({\bf{r}})\|$, where $\|{\bf{F}}_{k}({\bf{r}})\|=\sqrt{\left<{{\bf{F}}_{k}({\bf{r}})}\,,{{\bf{F}}_{k}({\bf{r}})}\right>}$. For simplicity, in the following, we consider the same structure for all elements so that $\|{\bf{F}}_{k}({\bf{r}})\|=\alpha,\forall k$. As a consequence, the relationship between ${\bf{j}}_{\mathrm{s}}$ and ${\bf{i}}_{\mathrm{s}}$ is ${\bf{j}}_{\mathrm{s}}=\alpha\,{\bf{i}}_{\mathrm{s}}$, and the total current density can be expressed as a function of the ports’ currents
${\mathsf{J}}({\bf{r}})=\sum_{k=1}^{K}{{\mathsf{J}}_{\mathrm{s}}}_{k}({\bf{r}})=\sum_{k=1}^{K}[{\bf{i}}_{\mathrm{s}}]_{k}\,\alpha\,{\mathbf{\Phi}}_{k}({\bf{r}})\,.$ (37)
@@
Exploiting the reciprocity theorem, the open circuit voltage at the $k$-th element’s port is $[{\bf{v}}]_{k}=\alpha\,\left<{{\mathsf{E}}({\bf{r}})}\,,{{\mathbf{\Phi}}_{k}({\bf{r}})}\right>=\alpha\,[{\bf{e}}]_{k}$.
In the case of a Hertzian dipole of infinitesimal length $l\ll\lambda$ located at position ${\bf{r}}_{k}$ with a generic polarization (orientation) $\hat{\bf{p}}_{k}$, the corresponding basis set is ${\mathbf{\Phi}}_{k}({\bf{r}})=\hat{\bf{p}}_{k}\,\delta({\bf{r}}-{\bf{r}}_{k})$. The current and open circuit voltage at the ports are, respectively, ${\bf{v}}=l\,{\bf{e}}$ and ${\bf{j}}_{\mathrm{s}}=l\,{\bf{i}}_{\mathrm{s}}$.
@@
Figure 8: The generic DSA structure with $N_{\text{A}}$ RF chains and $N_{\text{A}}$ active elements surrounded by $N_{\text{S}}$ reconfigurable scattering elements. Example of MIMO communication in which the DSA is adopted at the transmitter.
V-B EM Sigmal Processing in the Reactive Near Field
@@
We illustrate here two examples related to the use of the general framework in Sec. III to model and design a dynamic scattering array (DSA) and a stacked intelligent metasurface (SIM) whose aim is to realize a given signal processing task at the EM level and minimize the number $N_{\text{A}}$ of RF chains.
@@ V-B1 Dynamic Scattering Arrays
With reference to Fig. 8, a DSA is composed of $K$ antenna elements (e.g., dipoles) located within the space ${\cal{S}}$ of which $N_{\text{A}}$ are active and hence connected to an RF chain, and $N_{\text{S}}=K-N_{\text{A}}$ are passive scattering antenna elements [20, 29, 46, 30]. The response of the DSA can be changed by loading each scattering antenna element with reconfigurable loads of impedances $\mathbf{z}_{\text{L}}=\mathbf{z}_{\text{L}}(\bm{\theta})$. Ideally, the loads are designed to be only reactive to avoid power losses, i.e., $\mathbf{z}_{\text{L}}=\jmath\bm{\theta}$, with $\bm{\theta}\in\mathbb{R}^{K}$. Now, the $N_{\text{A}}$ RF chains introduce impressed open circuit voltages ${{\bf{v}}_{\mathrm{i}}=[v_{1},v_{2},\ldots,v_{N_{\text{A}}},0,0,\ldots,0]^{{\mathrm{T}}}}$ to the first $N_{\text{A}}$ elements that carry the information and are responsible for the impressed field ${\bf{e}}_{\mathrm{i}}$. In particular, in Sec. III we have seen that it should be ${\bf{e}}={\bf{e}}_{\mathrm{i}}+{\bf{e}}_{\mathrm{s}}$, where ${\bf{e}}={\bf{D}}^{-1}{\bf{j}}_{\mathrm{s}}$ and ${\bf{e}}_{\mathrm{s}}={\bf{G}}\,{\bf{j}}_{\mathrm{s}}$. Moreover, since $\{{\mathbf{\Phi}}_{k}({\bf{r}})\}$’s are disjoint functions, then ${\bf{D}}$ is diagonal. By multiplying the first relationship by the constant $\alpha$ defined in Sec. V-A, it is possible to map it into a relationship between currents and voltages at the ports, i.e., $\alpha^{2}\,{\bf{D}}^{-1}{\bf{i}}_{\mathrm{s}}={\bf{v}}_{\mathrm{i}}+\alpha^{2}{\bf{G}}\,{\bf{i}}_{\mathrm{s}}$. After simple steps, we obtain
$\left(\jmath{\rm diag}\left(\bm{\theta}\right)-{\bf{Z}}\right){\bf{i}}_{\mathrm{s}}={\bf{v}}_{\mathrm{i}}$ (38)
where ${\bf{Z}}=\alpha^{2}\,{\bf{G}}$ is the impedance matrix of the antenna structure and $\mathbf{z}_{\text{L}}(\bm{\theta})=\jmath{\rm diag}\left(\bm{\theta}\right)={\bf{D}}^{-1}$ represents the set of reconfigurable loads’ impedance. All the interactions between the elements of the DSA are captured by the impedance matrix ${\bf{Z}}$, which does not depend on the reconfigurable loads, and relates the voltages and currents of the $K$ ports [35]. With Hertzian dipoles, it can be computed analytically as $[{\bf{Z}}]_{n,k}=l^{2}\,\hat{\bf{p}}_{n}\bm{\cdot}\mathsf{G}\left({\bf{r}}_{n}-{\bf{r}}_{k}\right)\,\hat{\bf{p}}_{k}$, with $\mathsf{G}\left({\bf{r}}\right)$ given by (2).
@@
For a given receiver space, the end-to-end mapping between the open circuit voltage at the $N_{\text{A}}$ ports and the observed field ${\bf{y}}$ is
${\bf{y}}=\alpha\,\mathbf{G}_{\text{R}}{\bf{U}}\,\left(\jmath{\rm diag}\left(\bm{\theta}\right)-{\bf{Z}}\right)^{-1}{\bf{v}}_{\mathrm{i}}\,.$ (39)
@@
The previous equation can be used as a general model for the DSA in an optimization problem once the desired response ${\bf{y}}$ is fixed.
@@
One design approach is named characteristic mode analysis and consists of designing $\bm{\theta}$ such that the current ${\bf{i}}_{\mathrm{s}}$, needed to obtain the desired response ${\bf{y}}$, represents the dominant scattering mode of the structure. This can be simply accomplished by selecting $\bm{\theta}$ such that the current ${\bf{i}}_{\mathrm{s}}$ resonates, i.e., $[\bm{\theta}]_{k}=\frac{1}{[{\bf{i}}_{\mathrm{s}}]_{k}}\left[\Im\{{\bf{Z}}\}{\bf{i}}_{\mathrm{s}}\right]_{k}$, for $k=1,2,\ldots,K$, where $\Im\{x\}$ denotes the imaginary part of $x$. This approach, originally proposed in [19], has been widely exploited for the design of ESPARs with a single RF chain [29, 46]. Unfortunately, this simple approach is not accurate when applied to large structures in which secondary modes might deviate from the actual response from the desired one, and it is not applicable in the case of multiple active antennas.
@@
More in general, one might want to solve the following constrained optimization problem for a given objective end-to-end channel matrix ${\bf{H}_{\text{o}}}\in\mathbb{C}^{M\times N_{\text{A}}}$
$\hat{\bm{\theta}}=\arg\min_{\bm{\theta}}\left\|\,\alpha\,\mathbf{G}_{\text{R}}{\bf{U}}\,\left(\jmath\,{\rm diag}\left(\bm{\theta}\right)-{\bf{Z}}\right)^{-1}-{\bf{H}_{\text{o}}}\right\|_{\mathrm{F}}$ (40)
@@
constrained to a fixed radiated power $P_{\text{T}}={\bf{i}}_{\mathrm{s}}^{{\mathrm{H}}}\Re\{{\bf{Z}}\}\,{\bf{i}}_{\mathrm{s}}$ [30], where $\Re\{x\}$ denotes the real part of $x$. The $k$-th row of ${\bf{H}_{\text{o}}}$ represents the desired end-to-end channel response associated with the $n$-th port of the DSA. It is worth noticing that with the same configuration $\hat{\bm{\theta}}$ of the DSA, $N_{\text{A}}$ different responses are obtained simultaneously, each one associated with the input port of one specific RF chain.

### v3

@@ IV-C DoF of the Cascade Link including a Reconfigurable Scattering Device
$C=\sum\limits_{k=1}^{K}\log_{2}\left(1+\left(\tilde{\sigma}_{k}^{(\mathrm{T})}\right)^{2}\left(\tilde{\sigma}_{k}^{(\mathrm{R})}\right)^{2}\frac{p_{k}}{\sigma_{\mathrm{noise}}^{2}}\right)\,$, (36)
@@
where $p_{k}=\max\left(\mu-\frac{\sigma_{\mathrm{noise}}^{2}}{\left(\tilde{\sigma}_{k}^{(\mathrm{T})}\right)^{2}\left(\tilde{\sigma}_{k}^{(\mathrm{R})}\right)^{2}},0\right)$, with $\mu$ obtained by fulfilling the identity $\sum\nolimits_{k=1}^{K}{p_{k}}=P_{\text{T}}$, and being $\tilde{\sigma}_{k}^{(\mathrm{T})}$ and $\tilde{\sigma}_{k}^{(\mathrm{R})}$ the singular values of ${\mathbf{H}}_{\mathrm{T}}$ and ${\mathbf{H}}_{\mathrm{R}}$, respectively. [^7] We have $\tilde{\sigma}_{k}^{(\mathrm{T})}=\sigma_{k}^{(\mathrm{T})}$, since $\mathbf{V}$} is unitary. In contrast, $\tilde{\sigma}_{k}^{(\mathrm{R})}$ may differ from $\sigma_{k}^{(\mathrm{R})}$ due to the presence of the matrix ${\mathbf{G}}^{-1}$ in ${\mathbf{H}}_{\mathrm{R}}$, which accounts for the coupling between the current density and the scattered electric field at the reconfigurable scattering device. It is important to note that the last $K-{N_{\text{c}}}$ power coefficients are zero due to the maximum rank of $\mathbf{H}(\bm{\theta})$ being $N_{\text{c}}$}. Specifically, because of the diagonal structure of $\mathbf{G}_{\text{T}}$ and $\mathbf{G}_{\text{R}}$, where only the first $N$ and $M$ diagonal entries, respectively, correspond to significant singular values, only the first $N_{\text{c}}$} elements of the symbol vector $\mathbf{x}$} (out of $K$) carry “usable” information data. In addition, (36) is obtained by setting $\mathbf{Q}={\rm diag}\left(p_{1},p_{2},\ldots p_{K}\right)$ and the coupling matrix ${\mathbf{C}}={\mathbf{I}}_{K}$. Under the Born approximation, from (20), the optimal device matrix should satisfy the condition
${\mathbf{R}}(\bm{\theta})={\mathbf{V}}_{\text{R}}{\mathbf{U}}^{{\mathrm{H}}}_{\text{T}}\,$, (37)
@@
where ${\mathbf{V}}_{\text{R}}$ and ${\mathbf{U}}_{\text{T}}$ represent the left and right eigenvector matrices of the singular value decomposition (SVD) of ${\mathbf{H}}_{\mathrm{T}}$ and ${\mathbf{H}}_{\mathrm{R}}$, respectively. From (37), it can be inferred that the end-to-end capacity is maximized by designing ${\mathbf{R}}(\bm{\theta})$ (and consequently ${\mathbf{D}}(\bm{\theta})$) in a way that aligns with the left eigenvectors of the first link and the right eigenvectors of the second link. This ensures that the end-to-end transmitter-device-receiver channel is diagonalized. Furthermore, the water-filling power allocation is applied to the ordered product of the singular values of the individual transmitter-device and device-receiver channels.
@@
The physical realizability of the optimal device matrix $\mathbf{R}(\bm{\theta})$ depends on the specific technology being used. As mentioned in Section III, when the boundary conditions imposed by the scattering device are local, such as in conventional RISs, the matrix $\mathbf{R}(\bm{\theta})$ is constrained to be diagonal, meaning that (37) can only be approximated [43]. Recently, new RIS structures have been introduced, which enable reconfigurable coupling between the elements, leading to a non-diagonal matrix (non-diagonal RISs) [44].
V Enabling EM Technologies and SP Methods
@@
In this section, we first provide a connection between EM and circuit-based representations, then a comprehensive overview of key enabling EM technologies that can be utilized within the ESP framework by presenting specific examples of their application in practical processing tasks.
@@
V-A Port-controlled Reconfigurable Scattering Devices: Equivalent Circuit Model
In most practical implementations, the reconfigurable scattering device consists of a discrete set of $K$ antenna elements positioned at locations ${\mathbf{r}}_{k}$, whose scattering characteristics can be adjusted by varying the load at their ports. In this section, we illustrate how the general framework outlined in Section III can be linked to circuit theory that is typically adopted to model multiport systems. With a limited loss of generality, for the $k$-th antenna element of the reconfigurable scattering device, let ${\mathbf{F}}_{k}({\mathbf{r}})$ $1/\mathrm{m}^{2}]$ denote the corresponding current density normalized to an input port current of 1 Ampere. Here, we are implicitly assuming that the antenna element can be characterized by only one dominant mode, which is a good approximation, for instance, for Hertzian dipoles and resonant structures such as the half-wave dipole.
The function ${\mathbf{F}}_{k}({\mathbf{r}})$ depends on the particular structure of the element as well as its position ${\mathbf{r}}_{k}$ and orientation. Such a distribution can be obtained in general through EM-based simulations, but closed-form expressions are available for simple structures such as Hertzian or half-wave dipoles [35].
@@ V-A Port-controlled Reconfigurable Scattering Devices: Equivalent Circuit Model
The entire scattering device can be considered as a linear $K$-port network where we define the complex voltage and current envelopes (in the following denoted simply as voltages and currents) at the ports, respectively, $\mathbf{v}$} (open circuit voltage) and ${\mathbf{i}}_{\mathrm{s}}$. The actual current density of the $k$-th element is therefore ${{\mathsf{J}}_{\mathrm{s}}}_{k}({\mathbf{r}})=[{{\mathbf{i}}_{\mathrm{s}}}]_{k}\cdot{\mathbf{F}}_{k}({\mathbf{r}})$. The corresponding unitary-energy basis function is ${\mathbf{\Phi}}_{k}({\mathbf{r}})={\mathbf{F}}_{k}({\mathbf{r}})/\|{\mathbf{F}}_{k}({\mathbf{r}})\|$, where $\|{\mathbf{F}}_{k}({\mathbf{r}})\|=\sqrt{\left<{{\mathbf{F}}_{k}({\mathbf{r}})}\,,{{\mathbf{F}}_{k}({\mathbf{r}})}\right>}$. For simplicity, in the following, we consider the same structure for all elements so that $\|{\mathbf{F}}_{k}({\mathbf{r}})\|=\alpha,\forall k$. As a consequence, the relationship between ${\mathbf{j}}_{\mathrm{s}}$ and ${\mathbf{i}}_{\mathrm{s}}$ is ${\mathbf{j}}_{\mathrm{s}}=\alpha\,{\mathbf{i}}_{\mathrm{s}}$, defined in Section III-C, and the total current density can be expressed as a function of the ports’ currents
${\mathsf{J}}({\mathbf{r}})=\sum_{k=1}^{K}{{\mathsf{J}}_{\mathrm{s}}}_{k}({\mathbf{r}})=\sum_{k=1}^{K}[{\mathbf{i}}_{\mathrm{s}}]_{k}\,\alpha\,{\mathbf{\Phi}}_{k}({\mathbf{r}})\,.$ (38)
@@
Exploiting the reciprocity theorem, the open circuit voltage at the $k$-th element’s port is ${\mathbf{v}}]_{k}=\alpha\,\left<{{\mathsf{E}}({\mathbf{r}})}\,,{{\mathbf{\Phi}}_{k}({\mathbf{r}})}\right>=\alpha\,[{\mathbf{e}}]_{k}$.
In Section III-C, we have seen that the vectors representing the total, incident, and scattered EM fields are related by ${\mathbf{e}}={\mathbf{e}}_{\mathrm{i}}+{\mathbf{e}}_{\mathrm{s}}$, where ${\mathbf{e}}={\mathbf{D}}^{-1}{\mathbf{j}}_{\mathrm{s}}$ and ${\mathbf{e}}_{\mathrm{s}}={\mathbf{G}}\,{\mathbf{j}}_{\mathrm{s}}$. Moreover, since $\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\}$’s are disjoint functions, then $\mathbf{D}$} is diagonal. By multiplying the first relationship by the constant $\alpha$, it is possible to map it into a relationship between currents and voltages at the ports, i.e., $\alpha^{2}\,{\mathbf{D}}^{-1}{\mathbf{i}}_{\mathrm{s}}={\mathbf{v}}_{\mathrm{i}}+\alpha^{2}{\mathbf{G}}\,{\mathbf{i}}_{\mathrm{s}}$, being ${\mathbf{v}}_{\mathrm{i}}=[v_{1},v_{2},\ldots,v_{K}]^{{\mathrm{T}}}$} the impressed open circuit voltages associated with the incident field ${\mathbf{e}}_{\mathrm{i}}$. After simple steps, we obtain
$-\left(\mathbf{Z}_{\text{L}}(\bm{\theta})+{\mathbf{Z}}\right){\mathbf{i}}_{\mathrm{s}}={\mathbf{v}}_{\mathrm{i}}\,$, (39)
where ${\mathbf{Z}}=\alpha^{2}\,{\mathbf{G}}$ is the impedance matrix of the antenna structure and $\mathbf{Z}_{\text{L}}(\bm{\theta})=-\alpha^{2}\,{\mathbf{D}}^{-1}$ is a diagonal impedance matrix accounting for the boundary conditions depending on the loads and possibly mutual coupling between the elements. [^8] The sign change is a consequence of the convention followed in Fig. 7 for current direction. Equation (39) corresponds to the multiport circuit depicted in Fig. 7, and represents the equivalent voltage/current formulation of (17). In fact, after simple substitutions, it is
${\mathbf{e}}_{\mathrm{s}}=-{\mathbf{Z}}\left(\mathbf{Z}_{\text{L}}(\bm{\theta})+{\mathbf{Z}}\right)^{-1}\,{\mathbf{e}}_{\mathrm{i}}={\mathbf{R}}(\bm{\theta})\,{\mathbf{e}}_{\mathrm{i}}\,$, (40)
which gives the reflection matrix ${\mathbf{R}}(\bm{\theta})$ as a function of the circuit’s impedances. Therefore, Fig. 7 can be seen as the multiport circuit version of the scheme in Fig. 4 that involves EM quantities. All the couplings between the elements of the structure are captured by the impedance matrix $\mathbf{Z}$}, which does not depend on the reconfigurable loads and relates the open-circuit voltages and currents of the $K$ ports [35]. In the case Hertzian dipoles of infinitesimal length $l\ll\lambda$ located at position ${\mathbf{r}}_{k}$ with a generic polarization (orientation) $\hat{\mathbf{p}}_{k}$, it can be computed analytically as ${\mathbf{Z}}]_{n,k}=l^{2}\,\hat{\mathbf{p}}_{n}\bm{\cdot}\mathsf{G}\left({\mathbf{r}}_{n}-{\mathbf{r}}_{k}\right)\,\hat{\mathbf{p}}_{k}$, with $\mathsf{G}\left({\mathbf{r}}\right)$ given by (3). The real part of the impedance matrix $\mathbf{Z}$} affects the radiated EM field whose power is given by $P_{\text{T}}={\mathbf{i}}_{\mathrm{s}}^{{\mathrm{H}}}\Re\{{\mathbf{Z}}\}\,{\mathbf{i}}_{\mathrm{s}}$ for a lossless structure, where $\Re\{x\}$ denotes the real part of $x$.
In general, the equivalent multiport circuit is accurate provided that the impedance matrix $\mathbf{Z}$} is well characterized. This typically requires full-wave simulations or measurements, especially for complex structures. Thus, while the equivalent multiport circuit is highly convenient for signal processing and communication theorists, it does not by itself reveal the underlying physical phenomena, even though these are inherently embedded in the circuit model.
Figure 7: Equivalent multiport circuit of the reconfigurable scattering device in Fig. 4. The circles represent voltage sources.
@@
Figure 8: The generic DSA structure with $N_{\mathrm{A}}$ RF chains and $N_{\mathrm{A}}$ active elements surrounded by $N_{\text{S}}$ reconfigurable scattering elements. Example of MIMO communication in which the DSA is adopted at the transmitter. The table reports the actual end-to-end matrix $\hat{\mathbf{H}}_{\text{o}}$ obtained by optimizing (42).
V-B EM Signal Processing in the Reactive Near-Field
@@
We illustrate here two examples related to the use of the general framework illustrated in the previous sections to model and design a dynamic scattering array (DSA) and a stacked intelligent metasurface (SIM) whose aim is to realize a given signal processing task at the EM level and minimize the number $N_{\mathrm{A}}$ of RF chains.
@@ V-B1 Dynamic Scattering Arrays
With reference to Fig. 8, a DSA consists of $K$ antenna elements (e.g., dipoles) located within the space $\cal{S}$}, of which $N_{\mathrm{A}}$ are active and connected to an RF chain, while the remaining $N_{\text{S}}=K-N_{\mathrm{A}}$ are passive scattering elements [20, 25, 45, 26]. The response of the DSA can be changed by loading each scattering antenna element with reconfigurable loads of impedances $\mathbf{z}_{\text{L}}=\mathbf{z}_{\text{L}}(\bm{\theta})$. Ideally, the loads are designed to be only reactive to avoid power losses, i.e., $\mathbf{z}_{\text{L}}=\jmath\bm{\theta}$, with $\bm{\theta}\in\mathbb{R}^{K}$. Now, the $N_{\mathrm{A}}$ RF chains introduce impressed open circuit voltages ${\mathbf{v}}_{\mathrm{i}}=[v_{1},v_{2},\ldots,v_{N_{\mathrm{A}}},0,0,\ldots,0]^{{\mathrm{T}}}$} to the first $N_{\mathrm{A}}$ elements that carry the information and are responsible for the impressed field ${\mathbf{e}}_{\mathrm{i}}$. To account for this, define the identity matrix ${\mathbf{Q}}\in\mathbb{R}^{K\times N_{\mathrm{A}}}$ so ${\mathbf{v}}_{\mathrm{i}}={\mathbf{Q}}\,{\mathbf{v}}_{\mathrm{a}}$, with ${\mathbf{v}}_{\mathrm{a}}=[v_{1},v_{2},\ldots,v_{N_{\mathrm{A}}}]^{{\mathrm{T}}}$. The DSA is governed by the relationship (39) where the impedance matrix $\mathbf{Z}$} embeds also the $N_{\mathrm{A}}$ active sources in this case. To ensure the maximum power transfer between the sources (i.e., the RF chains) and the DSA, a power matching network has to be inserted, even though it is not explicitly shown in the figure for simplicity. Since the input impedance of the $N_{\mathrm{A}}$ ports of the DSA depends on the DSA configuration $\bm{\theta}$, the matching network should be configured accordingly [46].
@@
For a given receiver space, from (35) and (39), by setting ${\mathbf{H}}_{\mathrm{T}}={\mathbf{I}}_{K}$ since the source and scattering spaces coincides, the end-to-end mapping between the open circuit voltage ${\mathbf{v}}_{\mathrm{a}}$ at the $N_{\mathrm{A}}$ ports and the observed field $\mathbf{y}$} is
${\mathbf{y}}=-\alpha\,\mathbf{G}_{\text{R}}{\mathbf{U}}\,\left(\mathbf{Z}_{\text{L}}(\bm{\theta})+{\mathbf{Z}}\right)^{-1}{\mathbf{v}}_{\mathrm{i}}+{\mathbf{n}}=-\alpha\,\mathbf{G}_{\text{R}}{\mathbf{U}}\,\left(\mathbf{Z}_{\text{L}}(\bm{\theta})+{\mathbf{Z}}\right)^{-1}{\mathbf{Q}}\,{\mathbf{v}}_{\mathrm{a}}+{\mathbf{n}}\,.$ (41)
@@
The previous equation can be used as a general model for the DSA in an optimization problem once the desired response $\mathbf{y}$} is fixed.
@@
One design approach is named characteristic mode analysis and consists of designing $\bm{\theta}$ such that the current ${\mathbf{i}}_{\mathrm{s}}$, needed to obtain the desired response $\mathbf{y}$}, represents the dominant scattering mode of the structure. This can be simply accomplished by selecting $\bm{\theta}$ such that the current ${\mathbf{i}}_{\mathrm{s}}$ resonates, i.e., $\bm{\theta}]_{k}=\frac{1}{[{\mathbf{i}}_{\mathrm{s}}]_{k}}\left[\Im\{{\mathbf{Z}}\}{\mathbf{i}}_{\mathrm{s}}\right]_{k}$, for $k=1,2,\ldots,K$, where $\Im\{x\}$ denotes the imaginary part of $x$. This approach, originally proposed in [19], has been widely exploited for the design of ESPARs with a single RF chain [25, 45]. Unfortunately, this simple approach is not accurate when applied to large structures in which secondary modes might deviate from the actual response from the desired one, and it is not applicable in the case of multiple active antennas.
@@
More in general, one might want to solve the following constrained optimization problem of the type in (21) for a given objective end-to-end channel matrix ${\mathbf{H}_{\text{o}}}\in\mathbb{C}^{K\times N_{\mathrm{A}}}$
$\hat{\bm{\theta}}=\arg\min_{\bm{\theta}}\left\|\,-\alpha\,\mathbf{G}_{\text{R}}{\mathbf{U}}\,\left(\mathbf{Z}_{\text{L}}(\bm{\theta})+{\mathbf{Z}}\right)^{-1}{\mathbf{Q}}-{\mathbf{H}_{\text{o}}}\right\|_{\mathrm{F}}\,$, (42)
@@
constrained to a fixed radiated power $P_{\text{T}}$ [26]. The $n$-th column of $\mathbf{H}_{\text{o}}$} represents the desired end-to-end channel response associated with the $n$-th port of the DSA, for $n=1,2,\ldots,N_{\mathrm{A}}$. It is worth noticing that with the same configuration $\hat{\bm{\theta}}$ of the DSA, $N_{\mathrm{A}}$ different responses are obtained simultaneously, each one associated with the input port of one specific RF chain.